OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -4, 16, ... if there are 7 terms? Choose one answer. a. -5,461 b. 5,461 c. -3,277 d. 3,277

OpenStudy (anonymous):

Someone please helP!

jimthompson5910 (jim_thompson5910):

The first term is 1. So a = 1 The common ratio is -4. So r = -4 ----------------------------------- Plug these values into the formula Sn = a*(1-r^n)/(1-r) to compute the nth partial sum Sn = a*(1-r^n)/(1-r) Sn = 1*(1-(-4)^n)/(1-(-4)) Sn = 1*(1-(-4)^n)/5 S7 = 1*(1-(-4)^7)/5 S7 = 1*(1-(-16384))/5 S7 = 1*(1+16384)/5 S7 = 3227 So the 7th partial sum if 3227, which means that the sum of the first seven terms is 3227

jimthompson5910 (jim_thompson5910):

oops meant to write 3277 and not 3227...

jimthompson5910 (jim_thompson5910):

but hopefully you get the idea